A car travels starting from rest 20m/s in 4second. It is travels by this velocity decrease becomes stop in 10 seconds

Draw the velocity graph

Find the acceleration in each section

Find the total desplacment


1
Expert's answer
2021-04-05T11:07:45-0400
"a_1=\frac{20-0}{4}=5\frac{m}{s^2}\\
a_2=\frac{0-20}{10}=-2\frac{m}{s^2}\\s=0.5(20)(14)=140\ m"